# RP1M: A Large-Scale Motion Dataset for Piano Playing with Bi-Manual Dexterous Robot Hands

### v1

@@ 6 Limitations & Conclusion
Limitations   Our paper has limitations in several aspects. Firstly, although our method lifts the requirement of human-annotated fingering, enabling RL training on diverse songs, our method still fails to achieve strong performance on challenging songs due to fast rhythms and mechanical limitations of the robot hands. This could be solved by proposing a better RL method and improving the hardware design of the robot hands. Secondly, our dataset only includes proprioceptive observations. However, humans play piano with multi-modal inputs, including vision, tactile sensing, and auditory information. Enabling the agent to play the piano from such rich input sources is an intriguing direction. Lastly, although we demonstrate better zero-shot generalization performance than RoboPianist-MT [4], there is still a gap between our best multi-task agent and RL specialists, which requires future investigation.
Conclusion   In this paper, we propose a large-scale motion dataset named RP1M for piano playing with bi-manual dexterous robot hands. RP1M includes 1 million expert trajectories for playing 2k musical pieces. To collect such a diverse dataset for piano playing, we lift the need for human-annotated fingering in the previous method by introducing a novel automatic fingering annotation approach based on optimal transport. On single songs, our method matches the baselines with human-annotated fingering and can be adopted across different embodiments. Furthermore, we benchmark various imitation learning approaches for multi-song playing. We report new state-of-the-art results in motion synthesis for novel music pieces and identify the gap to achieving human-level piano-playing ability. We believe the RP1M dataset, with its scale and quality, forms a solid step towards empowering robots with human-level dexterity.
@@ Acknowledgments
We thank the support of the Max Planck Institute for Intelligent Systems, Tübingen (Germany). We acknowledge CSC – IT Center for Science, Finland, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through CSC. Yi Zhao, Juho Kannala, and Joni Pajarinen acknowledge funding by the Research Council of Finland (345521 353138, 327911). We thank Yuxin Hou and Wenyan Yang for the insightful discussion.
@@
Appendix A RP1M Dataset Collection Details
A.1 Reward formulation
@@ A.2 Training details
A.3 Computational resources
We train our RL agents on the cluster equipped with AMD MI250X GPUs, 64 cores AMD EPYC “Trento” CPUs, and 64 GBs DDR4 memory. Each agent takes  21 hours to train. The overall data collection cost is roughly 21 hours * 2089 agents = 43,869 GPU hours.
A.4 MuJoCo XLA Implementation
@@
Appendix B Multitask Benchmarking Details
@@
B.1 Baselines and hyperparameters
B.1.1 BC
Behavior Cloning (BC) [62] directly learns a policy by using supervised learning on observation-action pairs from expert demonstrations, which is one of the simplest methods to acquire robotic skills. Due to its straightforward approach and proven efficacy, BC is popular across multiple fields. The method employs a Multi-Layer Perceptron (MLP) as the policy network. Given expert trajectories, the policy network learns to replicate expert behavior by minimizing the Mean Squared Error (MSE) between predicted and actual expert actions. Despite its advantages, BC tends to perform poorly in generalizing to unseen states from the expert demonstrations. In our study, we evaluated three MLP models with varying hidden dimensions—256, 1024, and 4096. The first two models feature three layers, while the model with 4096 hidden dimensions is designed with six layers.
Table 5: BC
@@ B.1 Baselines and hyperparameters
Learning Rate 3e-4
Activation GELU
Training Steps 1M
@@
B.1.2 IBC
Implicit Behavioral Cloning (IBC) [63] adopts a novel angle on behavior cloning by reformulating supervised imitation learning as a conditional energy-based modeling problem. It trains an implicit policy represented by an energy function that is conditioned on both the action and observation, utilizing the InfoNCE loss [66]. This method demonstrates improved generalization over traditional BC. However, it encounters typical difficulties associated with training energy-based models, and the need for intensive action sampling and optimization at inference time, which may not scale well to high-dimensional action spaces.
Table 6: IBC
@@ B.1.2 IBC
Learning Rate Scheduler cosine
Training Steps 1M
Weight Decay 1e-6
Prediction Num of Iteration 5
Prediction Num of Sample 1024
Observation Horizon 2
Prediction Horizon 2
@@
B.1.3 BC-RNN
BC-RNN [64] is a variant of BC that incorporates a Recurrent Neural Network as the policy network to capture a sequence of past observations. It is the best-performing baseline in the Robomimic paper [64].
Table 7: BC-RNN
@@ B.1.3 BC-RNN
Learning Rate Scheduler linear
Training Steps 1M
@@
Action Horizon 1
B.1.4 Diffusion Policy
Diffusion models have achieved many state-of-the-art results across image, video, and 3D content generation [67, 68, 69, 70, 71]. In the context of robotics, diffusion models have been used as policy networks for imitation learning in both manipulation [5, 72, 73, 74] and locomotion tasks [75], showing remarkable performance across various robotic tasks. Diffusion Policy [5] proposed to learn an imitation learning policy with a conditional diffusion model. It models the action distribution by inverting a process that gradually adds noise to a sampled action sequence, conditioning on a state and a sampled noise vector. We used a CNN-based Diffusion Policy with DDIM [65] as the sampler to diffuse out action trajectories for improved efficiency. We build our diffusion policy training pipeline based on the Robomimic [64] and DROID [76], which provide high-quality implementations.
Table 8: Diffusion Policy
@@ B.1.4 Diffusion Policy
Batch Size 128
Optimizer Adam
Learning Rate 1e-4
Learning Rate Scheduler Linear
Training Steps 1M
Diffusion Method DDIM
@@
U-Net Hidden Layer Sizes [256, 512, 1024]
Observation Horizon 2
@@
Action Horizon 1
B.2 Training and evaluation
We train the policies with 3 different sizes of expert data: 50, 150, and 300 songs, respectively. Subsequently, we assess the trained policies using three distinct categories of musical pieces. The first category, in-distribution songs, includes pieces that are part of the training datasets. Evaluating with in-distribution songs tests the multitasking abilities of the policies and checks if a policy can accurately recall the songs on which it was trained. The second group of songs for evaluation are easy out-of-distribution (OOD) songs: those music pieces do not overlap with the training songs but they are easy to play. They only contain slow motions and short horizons. The third group of evaluation songs are hard out-of-distribution songs: those are difficult music pieces that do not overlap with the training songs. They contain more diverse motions and longer horizons. This out-of-distribution evaluation measures the zero-shot generalization capabilities of the policies. Analogous to an experienced human pianist who can play new pieces at first sight, we aim to determine if it is feasible to develop a generalist agent capable of playing the piano under various conditions.
@@ B.2 Training and evaluation
Table 10: Easy out-of-distribution songs
RoboPianist-debug-TwinkleTwinkleLittleStar-v0
RoboPianist-debug-CMajorChordProgressionTwoHands-v0
RoboPianist-debug-TwinkleTwinkleRousseau-v0
RoboPianist-debug-NocturneRousseau-v0
RoboPianist-debug-NocturneRousseau-v0
Table 11: Hard out-of-distribution songs
GP-AkimenkoTheodoreAuCoinDuFeuOp28-v0
GP-AgnewRoy2PianoPieces-v0
GP-AlbaAntonioElEnsuenoOp16-v0
GP-AlbaAntonioSensitiva-v0
GP-MinotAdolfMisterioso-v0

### v2

@@ 6 Limitations & Conclusion
Limitations  Our paper has limitations in several aspects. Firstly, although our method lifts the requirement of human-annotated fingering, enabling RL training on diverse songs, our method still fails to achieve strong performance on challenging songs due to fast rhythms and mechanical limitations of the robot hands. Improving the RL method and hardware design could help address this. Secondly, the evaluation metric, F1 score, may not adequately capture musical performance and the position-based controller missing the target velocity would hinder the performance. Thirdly, our dataset includes only proprioceptive observations, whereas humans play piano using multi-modal inputs like vision, touch, and hearing; incorporating these could enhance the agent’s capabilities. Furthermore, there are several challenges to deploying the learned agent on a real-world robot. This includes the challenges of obtaining the state of the piano and the hands (e.g., tracking the precise fingertip positions), optimizing a precise position controller at high speed as well as the sim-to-real gap for the highly dynamic piano-playing task, etc. Lastly, although we demonstrate better zero-shot generalization performance than RoboPianist-MT [4], there is still a gap between our best multi-task agent and RL specialists, which requires future investigation.
Conclusion  In this paper, we propose a large-scale motion dataset named RP1M for piano playing with bimanual dexterous robot hands. RP1M includes 1 million expert trajectories for playing 2k musical pieces. To collect such a diverse dataset for piano playing, we lift the need for human-annotated fingering in the previous method by introducing a novel automatic fingering annotation approach based on optimal transport. On single songs, our method matches the baselines with human-annotated fingering and can be adopted across different embodiments. Furthermore, we benchmark various imitation learning approaches for multi-song playing. We report promising results in motion synthesis for novel music pieces when increasing the data size and identify the gap to achieving human-level piano-playing ability. We believe the RP1M dataset, with its scale and quality, forms a solid step towards empowering robots with human-level dexterity.
@@ Acknowledgments
We thank the support of the Max Planck Institute for Intelligent Systems, Tübingen (Germany). We acknowledge CSC – IT Center for Science, Finland, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through CSC. Yi Zhao, Juho Kannala, and Joni Pajarinen acknowledge funding by the Research Council of Finland (345521, 353138, 327911). We thank Yuxin Hou and Wenyan Yang for the insightful discussion. We thank all reviewers for their detailed and constructive comments.
@@
Appendix A More Related Work
Due to the difficulty of actually modeling the dynamics of the dexterous robot hand, recent methods resort to learning-based approaches, especially RL, which has achieved huge success in both robotics [57, 58, 1] and computer graphics [59]. To ease the training of dexterous robot hands with a large number of degrees of freedom (DoFs), demonstrations are commonly used [60, 61, 62, 63, 64]. Due to the advance of both RL algorithms and simulation, recent work shows impressive results on dexterous hand manipulation tasks without human demonstrations. Furthermore, the policy trained in the simulator can further be deployed on real dexterous robot hands via sim-to-real transfer [65, 66, 67, 30, 68, 69, 70].
Generalist Agents RL methods usually perform well on single tasks, however, as human beings, we can perform multiple tasks. Generalist agents are proposed to master a diverse set of tasks with a single agent [71, 57, 72, 73]. These methods typically resort to scalable models and large datasets [72, 74, 75, 76, 77]. Recently, diffusion models have achieved many state-of-the-art results across image, video, and 3D content generation [78, 79, 80, 81, 82] In the context of robotics, diffusion models have been used as policy networks for imitation learning in both manipulation [5, 75, 83, 84] and locomotion tasks [85]. The same technique has also been investigated in multi-task learning [75, 84]. We investigate the application of diffusion policy in high-dimensional control tasks, that is, playing piano with bimanual dexterous robot hands.
Appendix B RP1M Dataset Collection Details
B.1 Reward Formulation
@@ B.2 Training Details
B.3 Computational Resources
We train our RL agents on the LUMI cluster equipped with AMD MI250X GPUs, 64 cores AMD EPYC “Trento” CPUs, and 64 GBs DDR4 memory. Each agent takes  21 hours to train. The overall data collection cost is roughly 21 hours * 2089 agents = 43,869 GPU hours.
B.4 MuJoCo XLA Implementation
@@
Appendix C Multitask Benchmarking Details
@@
C.1 Baselines and Hyperparameters
C.1.1 Behavior Cloning
Behavior Cloning (BC) [43] directly learns a policy using supervised learning on observation-action pairs from expert demonstrations, one of the simplest methods to acquire robotic skills. Due to its straightforward approach and proven efficacy, BC is popular across multiple fields. The method employs a Multi-Layer Perceptron (MLP) as the policy network. Given expert trajectories, the policy network learns to replicate expert behavior by minimizing the Mean Squared Error (MSE) between predicted and actual expert actions. Despite its advantages, BC tends to perform poorly in generalizing to unseen states from the expert demonstrations. The MLP we used features three hidden layers, each with 512 units, followed by Layer Normalization and an Exponential Linear Unit (ELU) activation function to stabilize training and introduce non-linearity.
Table 5: Hypermeters used in BC
@@ C.1.1 Behavior Cloning
Learning Rate 1e-4
@@
C.1.2 BeT
Behavior Transformers (BeT) [44] uses a transformer-decoder based backbone with a discrete action mode predictor coupled with a continuous action offset corrector to model continuous actions sequences. It clusters continuous actions into discrete bins using k-means to model high-dimensional, continuous multi-modal action distributions as categorical distributions without learning complicated generative models. We adopted the implementation and hyperparameters from the Diffusion Policy codebase [5].
Table 6: Hyperparamerters used in BeT
@@ C.1.2 BeT
Num of bins 64
MinGPT n_layer 8
MinGPT n_head 8
MinGPT n_embd 120
Observation Horizon 1
Prediction Horizon 1
@@
C.1.3 Diffusion Policy
Diffusion models have achieved many state-of-the-art results across image, video, and 3D content generation [78, 79, 80, 81, 82]. In the context of robotics, diffusion models have been used as policy networks for imitation learning in both manipulation [5, 75, 83, 84] and locomotion tasks [85], showing remarkable performance across various robotic tasks. Diffusion Policy [5] proposed to learn an imitation learning policy with a conditional diffusion model. It models the action distribution by inverting a process that gradually adds noise to a sampled action sequence, conditioning on a state and a sampled noise vector. We evaluated both the U-Net-based Diffusion Policy (DP-U) and the transformer-based Diffusion Policy (DP-T). We build our diffusion policy training pipeline based on the original Diffusion Policy [5] codebase, which provides high-quality implementations.
Table 7: Hyperparameters used in DP-U
@@ C.1.3 Diffusion Policy
Weight Decay 1e-6
Diffusion Method DDPM
Number of Diffusion Iterations 100
EMA Power 0.75
U-Net Hidden Layer Sizes [256, 512, 1024]
Diffusion Step Embedding Dim. 256
@@
Action Horizon 4
Table 8: Hyperparameters used in DP-T
@@
Batch Size 1024
Optimizer AdamW
Learning Rate 1e-3
Weight Decay 1e-4
Diffusion Method DDPM
@@
n_layer 8
n_head 4
n_emb 156
p_drop_emb 0.0
p_drop_attn 0.3
Observation Horizon 1
@@
Action Horizon 4
C.2 Training and Evaluation
We train the policies with 5 different sizes of expert data: 12, 25, 50, 100, and 150 songs, respectively. Subsequently, we assess the trained policies using two distinct categories of musical pieces. The first category, in-distribution songs, includes pieces that are part of the training datasets. Evaluating with in-distribution songs tests the multitasking abilities of the policies and checks if a policy can accurately recall the songs on which it was trained. The second group of songs for evaluation are out-of-distribution songs: those music pieces do not overlap with the training songs. The selected songs contain diverse motions and long horizons, making them challenging to play. This out-of-distribution evaluation measures the zero-shot generalization capabilities of the policies. Analogous to an experienced human pianist who can play new pieces at first sight, we aim to determine if it is feasible to develop a generalist agent capable of playing the piano under various conditions.
@@ C.2 Training and Evaluation
Table 10: Out-of-distribution songs
GiantMIDI-IsmagilovTimurSpringSketches-v0
GiantMIDI-JohnsonCharlesLeslieGoldenSpiderRag-v0
GiantMIDI-JoseffyRafaelValseDesDames-v0
GiantMIDI-KiefertCarlPastoral-v0
GiantMIDI-KleberHenryTheFancyPolka-v0
GiantMIDI-KockKarlMelancholie-v0
GiantMIDI-LackTheodoreMenuetDuXviiimeSiecleOp36-v0
GiantMIDI-LecocqCharlesLeJourEtLaNuit-v0
GiantMIDI-LefebureWelyLouisJamesAlfredApresLaChass-v0
GiantMIDI-LindgreenCharlesFirstCourtship-v0
GiantMIDI-LisztFranzRomanceOublieeS527-v0
GiantMIDI-MacchiClaudioRomanzaSenzaParole1-v0
GiantMIDI-MarcouPaulLeCosaqueOp15-v0
GiantMIDI-MasonWilliamAPastoralNovelette-v0
GiantMIDI-MattheyUlissePensieroOstinato-v0
GiantMIDI-MayerlBillyJosephEgyptianSuite-v0
GiantMIDI-MendelssohnFelixScherzoWoo2-v0
GiantMIDI-MozartWolfgangAmadeusAllegroInCMajorK484-v0
GiantMIDI-NegriCesareAlemanaDamore-v0
GiantMIDI-OkellyJosephEn1795Op52-v0